A college student earned $5100 during summer break. He invested all the money. Part of the money he invested at 8% and the rest at 6%. He earned a total of $362.00 in interest. How much did he invest at 8% and how much did he invest at 6%?
Answer by stanbon(75887)

A college student earned $5100 during summer break. He invested all the money. Part of the money he invested at 8% and the rest at 6%. He earned a total of $362.00 in interest. How much did he invest at 8% and how much did he invest at 6%?
-----
Quantity Eq: x + y = 5100
Interest Eq:0.08x + 0.06y = 362
------------------------
Multiply thru the 1st by 8
Multiply thru the 2nd by 100
8x + 8y = 8*5100
8x + 6y = 36200
-------------------
Subtract the 2nd from the 1st and solve for "y":
2y = 4600
y = $2300 (amt. invested at 6%
---
Since x+y = 5100, x = 5100-2300 = $2800 (amt. invested at 8%
=================
Cheers,
Stan H.